OpenStudy (anonymous):

Find the values of the 30th and 90th percentiles of the data;129,113,200,100,105,132,100,176

OpenStudy (anonymous):

@campbell_st ur the only 1 left homie

OpenStudy (campbell_st):

well I'd arrange the score in order, lowest to highest percentiles are 1/10 ths so 30% = 3/10 * 8 = 2.4 find the score in the 2.4th position for 90% find the score in the 7.2 position this is best guess... as I can't ever remember doing percentiles

OpenStudy (campbell_st):

best guess is 105-100 = 5 * 0.4 = 2 so the 30% percentile is 102 200-176 = 24 24*0.2 = 4.8 90% percentile 180.8